# Markov Perfect Equilibria in

### v1

@@ A.1 Dynamic programming for Markov chains
For $h>0$ and $K\geq 1$ let $\bar{\mathcal{T}}:=\{0,h,\ldots,Kh\}$ and ${\mathcal{T}}:=\bar{\mathcal{T}}\setminus\{Kh\}$. Consider a state process $(X_{t},Y_{t})_{t\in\bar{\mathcal{T}}}$ taking values in a discrete set ${\mathcal{X}}\times{\mathcal{Y}}$ where $|{\mathcal{X}}|=d_{1}<\infty$ and $|{\mathcal{Y}}|=d_{2}<\infty$. Fix a map $\beta:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{2}-1}$. For any control $\alpha:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{1}-1}$, let $\mathbb{P}^{\alpha}$ denote the probability measure under which $(X,Y)$ is a Markov chain with $\mathbb{P}^{\alpha}$-transition probabilities
@@
Let $w:\bar{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\mathbb{R}$ be given and assume the cost functions $\ell,g$ are continuous. Then $w=v$ if and only if $w$ solves the following dynamic programming equation, for $(t,x,y)\in\bar{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}$,
@@
Moreover, a control $\alpha:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{1}-1}$ is optimal for any initial data $(t,x,y)$ if and only if
@@
There are no issues with measurability since ${\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}$ is a discrete set. When the dynamics of $Y$ are deterministic, we may allow ${\mathcal{Y}}$ to be any set without changing the validity of Lemma A.1.
@@
Finally, the normal cone of the set $C$ is denote by $N_{C}$:
@@
Proof.
Step 1. We first show that $\eta:=f+\delta_{C}$ is a closed, proper and strongly convex function. Clearly, $\eta$ is proper. Closedness of $\eta$ is obvious by closedness of $f$, for example using closedness of the sublevel sets [7, Theorem 2.6]. Next, using $\partial\delta_{C}=N_{C}$, see [7, Example 3.5], and the the sum rule of subdifferential calculus, see [7, Theorem 3.40], we have
@@
where the right side is understood in the sense of Minkowski. Hence $\mathrm{dom}(\partial\eta)=\mathrm{dom}(\partial f)\cap\mathrm{dom}(\partial\delta_{C})=\mathrm{dom}(\partial f)\cap C$. Strong convexity means that for $(x,x^{\prime})\in\mathrm{dom}(\partial\eta)\times\mathrm{dom}(\eta)$ and $\phi\in\partial\eta(x)$,
$\eta(x^{\prime})\geq\eta(x)+\phi\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{2}^{2},$ (A.1)
see [7, Theorem 5.24]. Let such $(x,x^{\prime},\phi)$ be given. By the above observation, there exists $g\in\partial f(x)$ and $n\in N_{C}(x)$ with $\phi=g+n$. Since $x,x^{\prime}\in C$, by (3.1),
$\eta(x^{\prime})=f(x^{\prime})$ $\geq f(x)+g\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{p}|x^{\prime}-x|_{q}$
$\geq f(x)+g\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{2}^{2}=\eta(x)+g\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{2}^{2}$
@@
$\eta(x^{\prime})\geq\eta(x)+g\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{2}^{2}\geq\eta(x)+\phi\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{2}^{2},$
@@
$f(v_{2})\geq f(v_{1})+g_{1}\cdot w+\gamma|w|_{r}|w|_{r^{\prime}}\quad\text{and}\quad f(v_{1})\geq f(v_{2})-g_{2}\cdot w+\gamma|w|_{r}|w|_{r^{\prime}}.$
@@
$f(v_{2})\geq f(v_{1})+y_{1}\cdot w+\gamma|w|_{r}|w|_{r^{\prime}}\quad\text{and}\quad f(v_{1})\geq f(v_{2})-y_{2}\cdot w+\gamma|w|_{r}|w|_{r^{\prime}}.$
@@
which, using $w\neq 0$, implies (A.2). ∎

### v2

@@ A.1 Dynamic programming for Markov chains
For $h>0$ and $K\geq 1$, let $\bar{\mathcal{T}}:=\{0,h,\ldots,Kh\}$ and ${\mathcal{T}}:=\bar{\mathcal{T}}\setminus\{Kh\}$. Consider a state process $(X_{t},Y_{t})_{t\in\bar{\mathcal{T}}}$ taking values in a discrete set ${\mathcal{X}}\times{\mathcal{Y}}$, where $|{\mathcal{X}}|=d_{1}<\infty$ and $|{\mathcal{Y}}|=d_{2}<\infty$. Fix a map $\beta:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{2}-1}$. For any control $\alpha:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{1}-1}$, let $\mathbb{P}^{\alpha}$ denote the probability measure under which $(X,Y)$ is a Markov chain with $\mathbb{P}^{\alpha}$-transition probabilities
@@
Let $w:\bar{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\mathbb{R}$ be given and assume that the cost functions $\ell,g$ are continuous. Then $w=v$ if and only if $w$ solves the following dynamic programming equation, for $(t,x,y)\in{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}$,
@@
Moreover, a control $\alpha:{\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}\to\Sigma^{d_{1}-1}$ is optimal for any initial point $(t,x,y)$ if and only if
@@
There are no issues with measurability since ${\mathcal{T}}\times{\mathcal{X}}\times{\mathcal{Y}}$ is a discrete set. Moreover, when the dynamics of $Y$ are deterministic, we may allow $\mathcal{Y}$} to be any set without changing the validity of Lemma A.1.
@@
Finally, the normal cone of a set $C$ is denoted by
@@
Step 1. We first show that $\eta:=f+\delta_{C}$ is a closed, proper and strongly convex function. Clearly, $\eta$ is proper. Closedness of $\eta$ follows from the closedness of $f$, for example using closedness of the sublevel sets [7, Theorem 2.6]. Next, using $\partial\delta_{C}=N_{C}$, see [7, Example 3.5], and the sum rule of subdifferential calculus, see [7, Theorem 3.40], we have
@@
where the right side is understood in the sense of Minkowski. Hence $\mathrm{dom}(\partial\eta)=\mathrm{dom}(\partial f)\cap\mathrm{dom}(\partial\delta_{C})=\mathrm{dom}(\partial f)\cap C$. Strong convexity of $\eta$ means that, for $(x,x^{\prime})\in\mathrm{dom}(\partial\eta)\times\mathrm{dom}(\eta)$ and $\phi\in\partial\eta(x)$,
$\eta(x^{\prime})\geq\eta(x)+\phi\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{2}^{2},$ (A.1)
see [7, Theorem 5.24]. Let such $(x,x^{\prime},\phi)$ be given. By the above observation, there exists a $g\in\partial f(x)$ and $n\in N_{C}(x)$ with $\phi=g+n$. Since $x,x^{\prime}\in C$, by (3.1),
$\eta(x^{\prime})=f(x^{\prime})$ $\geq f(x)+g\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{p}|x^{\prime}-x|_{q}$
$\geq f(x)+g\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{2}^{2}=\eta(x)+g\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{2}^{2}$
@@
$\eta(x^{\prime})\geq\eta(x)+g\cdot(x^{\prime}-x)+n\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{2}^{2}=\eta(x)+\phi\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{2}^{2},$
@@
$f(v_{2})\geq f(v_{1})+g_{1}\cdot w+\frac{\gamma}{2}|w|_{r}|w|_{r^{\prime}}\quad\text{and}\quad f(v_{1})\geq f(v_{2})-g_{2}\cdot w+\frac{\gamma}{2}|w|_{r}|w|_{r^{\prime}}.$
@@
$f(v_{2})\geq f(v_{1})+y_{1}\cdot w+\frac{\gamma}{2}|w|_{r}|w|_{r^{\prime}}\quad\text{and}\quad f(v_{1})\geq f(v_{2})-y_{2}\cdot w+\frac{\gamma}{2}|w|_{r}|w|_{r^{\prime}}.$
@@
which, using $w\neq 0$, implies (A.2). $\Box$